OpenStudy (anonymous):

Calc AB Question The function v(t) = t^2-3t-10 is the velocity in m/sec for a particle moving along the x-axis, where t is measured in seconds (t >/= 0). a. Determine when the particle is moving to the right, to the left, and stopped on the interval 0

OpenStudy (praxer):

I think for the b. we need to do \[\int\limits_{0}^{7}v(t)\] @Kainui am i correct ?

OpenStudy (kainui):

Yeah, but remember the dt! Remember, distance=rate*time so literally an infinitesimal bit of movement will be: dx=v(t)*dt add up all the infinite number of infinitesimally small distances and you get:\[distance = \int\limits_{0}^{7}v(t)dt\]

OpenStudy (praxer):

:) I missed the dt.

OpenStudy (kainui):

All good, I just wanted to explain sort of the reasoning behind it to @MathGeekinProgress

OpenStudy (whpalmer4):

For part a) take the derivative of v(t), set it equal to 0 and solve for t to see where the velocity changes sign.

OpenStudy (anonymous):

Oh I know about the dt but honestly, where do I start? What about part a?

OpenStudy (anonymous):

Thanks @whpalmer4 for part a haha

OpenStudy (anonymous):

The derivative ov v(t) is 2t-t

OpenStudy (whpalmer4):

Check your work...

OpenStudy (kainui):

Well what makes v(t) different when it's moving to the left vs the right? How about when it's stopped?

OpenStudy (anonymous):

Wouldn't that be the same as a(t)?

OpenStudy (anonymous):

Woops. 2t -3*

OpenStudy (kainui):

The derivative of v(t) is a(t), yes.

OpenStudy (whpalmer4):

you know what, it's late here, scratch what I said...

OpenStudy (anonymous):

So 2t-3=0 is t=3/2

OpenStudy (anonymous):

What???

OpenStudy (whpalmer4):

No warranty of correctness applies :-)

OpenStudy (whpalmer4):

You're in good hands with these guys, I'm going to bed!

OpenStudy (anonymous):

How do I know where the sign changes?

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

a) v(t) < 0, the particle moves left v(t) > 0, the particle moves right b) ∫v(t) dt, from 0 to 7 c) remember, distance isn't the same as displacement. distance = ∫│v(t) │dt, from 0 to 7 I will let you do theh math

OpenStudy (kainui):

@MathGeekinProgress Yeah sourwing's on the right track.

OpenStudy (anonymous):

For part a, if v(t) = 0, then we know the particle is stopped, correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

So for part a, the particle is moving towards the riight, correct?

OpenStudy (anonymous):

v(t) = 0 = t^2-3t-10 = (t - 5) (t + 2), t = 5 (exclude - 2) so you have to check the interval (0,5) and (5,7) for the sign of v(t)

OpenStudy (anonymous):

I dont get it..

OpenStudy (anonymous):

For Part B I got -174/6. For Part C I got 1547/3.

OpenStudy (anonymous):

well, at t = 5 is when the particle stopped. Everywhere else is the particle either moving left of right. Like I said before, v(t) > 0 means the particle moves to the right and v(t) < 0 means the particle moves left

OpenStudy (anonymous):

B) it's -175/6 (typo?) C) it's 125/2

OpenStudy (anonymous):

Yeah typo

OpenStudy (anonymous):

How did you get 125/2?

OpenStudy (anonymous):

how did you get 1547/3?

OpenStudy (anonymous):

I did the same exact thing for part C as I did for part B. The only difference was that for part B I used t^2-3t-10. For part C I used t^2+3t+10

OpenStudy (anonymous):

|t^2 - 3t - 1| isn't the same as t^3 + 3t + 10

OpenStudy (anonymous):

t^2 - 3t - 10 if t^2 - 3t - 10 > 0 |t^2 - 3t - 10| = -(t^2 - 3t - 10) if t^2 - 3t - 10 <= 0

OpenStudy (anonymous):

Ohhhhhh. So it's gonna be - t^2 + 3t + 10 because v(t) = 0

OpenStudy (anonymous):

in other words: t^2 - 3t - 10 if 5 < t < 7 |t^2 - 3t - 10| = -(t^2 - 3t - 10) if 0 < t < 5

OpenStudy (anonymous):

\[distance = \int\limits_{0}^{5}-(t^2-3t-10)dt + \int\limits_{5}^{7} (t^2-3t-10)dt\]

OpenStudy (anonymous):

Ohhhhhhh

OpenStudy (anonymous):

let me solve this

OpenStudy (anonymous):

which turns out to be 275/6 + 50/3 = 125/2

OpenStudy (anonymous):

Yup I got that!

OpenStudy (anonymous):

Now what about part a?

OpenStudy (anonymous):

well, where is v(t) < 0? v(t) > 0?

OpenStudy (anonymous):